OpenStudy (anonymous):

evaluate the limit lim cos(7x) -1 / x^2 x->0

OpenStudy (anonymous):

so far i've done -7 sin 7x -1 / 2x

OpenStudy (anonymous):

using the double angle formula, we have \(cos 2\theta=1-2\sin^2\theta\)

OpenStudy (anonymous):

need to use l hospital's rule to evaluate but i'm not sure how to work it out

OpenStudy (anonymous):

no need for that.. just use the double angle formula

OpenStudy (anonymous):

\[ L=\lim_{x\to0}\frac{1-2\sin^2\left(7x\over2\right)-1}{x^2} \]

OpenStudy (anonymous):

follow?

OpenStudy (anonymous):

no I ended up with -7cos(7x) / 2 = 7/2

OpenStudy (anonymous):

our whole point was to get the sine

OpenStudy (anonymous):

use the equation editor below to enter your steps.

OpenStudy (anonymous):

the answer is 49/2

OpenStudy (anonymous):

i mean -49/2

OpenStudy (anonymous):

we are trying to create the form \[ \lim_{x\to0}\frac{\sin x}{x}=1 \]

OpenStudy (anonymous):

Thank you!